# A comprehensive survey on deep active learning and its applications in medical image analysis

### v1

@@ 6.1 Towards Active Learning with Better Uncertainty
Recently, many works have adopted learnable performance estimation for quality control of deep model outputs. For instance, the recently proposed segment anything model (SAM) (Kirillov et al., 2023) provides IoU estimates for each mask to evaluate its quality. In medical image analysis, automated quality control is critical to ensure the reliability and safety of the deep model outputs (Kohlberger et al., 2012). For example, Wang et al. (2020c) employed deep generative models for learnable quality control in cardiac MRI segmentation, where the predicted Dice scores showed a strong linear relationship with the real ones. Additionally, Billot et al. (2023) used an additional neural network to predict the Dice coefficient of brain tissue segmentation results. Overall, learnable performance estimation can accurately predict the quality of model outputs. Hence, delving deeper into their potential for uncertainty-based AL is crucial to effectively tackle the issue of over-confidence.
@@ 6.2 Towards Active Learning with Better Representativeness
Representativeness-based AL effectively utilizes feature representations and data distributions for sample selection. Cover-based and diversity-based AL methods implicitly capture the data distribution, whereas density-based AL explicitly estimates it. However, the latter requires supplementary strategies to ensure diversity. As the core of density-based AL, density estimation in high-dimensional spaces has always been challenging. Popular density estimation methods, such as kernel density estimation and GMM, may face challenges in high-dimensional spaces. In future research, we can consider introducing density estimators tailored to high-dimensional spaces.
@@ 6.3 Towards Active Learning with Weak Annotation
In §4.4, we discuss region-based active learning, which only requires region-level annotation of a sample. However, annotating all pixels within the region is still needed. Some AL works have incorporated weak annotations to simplify the task for annotators. In object detection tasks, Vo et al. (2022) trained deep models with image-level annotation. They selected samples with box-in-box prediction results and annotated them with bounding boxes. Moreover, Lyu et al. (2023) adopted disagreement to choose which objects are worth annotating. Rather than annotating all objects within the image, they only required box-level annotations for a subset of objects. In AL of instance segmentation, Tang et al. (2022) only required annotations for each object’s class label and bounding box, without the annotation of fine-grained segmentation masks. In future research, AL based on weak annotations is a direction worthy of in-depth exploration.
@@ 6.4 Towards Active Learning with Better Generative Models
In §4.5, we summarize the applications of generative models in AL. However, existing works have mainly focused on using GANs as sample generators. Recently, diffusion models (Kazerouni et al., 2023) have advanced in achieving state-of-the-art generative quality. Furthermore, text-to-image diffusion models, represented by Stable Diffusion (Rombach et al., 2022), have revolutionized the image generation domain. Their high-quality, text-guided generation results enable a more flexible image generation. Exploring the potential of diffusion models in deep AL is a promising avenue for future research.
@@ 6.5 Towards Active Learning with Foundation Models
With the rise of visual foundational models, such as contrastive language-image pretraining (CLIP) (Radford et al., 2021) and SAM (Kirillov et al., 2023), and large language models (LLMs) like GPT (OpenAI, 2023), deep learning in medical image analysis and computer vision is undergoing a paradigm shift. These foundational models (Bommasani et al., 2021) offer new opportunities for the development of AL.
AL is closely related to the training paradigms in deep learning of computer vision and medical image analysis. From the initial approach of train-from-scratch to the “pretrain-finetune” strategy using supervised or self-supervised pre-trained models, these paradigms usually require fine-tuning the entire network. Foundation models contain a wealth of knowledge. When combined with recently emerging parameter-efficient fine tuning (PEFT) or prompt tuning techniques (Hu et al., 2021; Jia et al., 2022), we can tune only a minimal subset of model weights (for example, 5%) for rapid transfer to downstream tasks. As the number of fine-tuned parameters decreases, AL has the potential to further reduce the number of required annotated samples. Therefore, it is essential to investigate the applicability of existing AL under PEFT or prompt tuning and explore the most suitable AL strategies for PEFT.

### v3

@@ 7.1 Towards Active Learning with Better Uncertainty
Recently, many works have adopted learnable performance estimation for quality control of deep model outputs. For instance, the recently proposed segment anything model (SAM) (Kirillov et al., 2023) provides IoU estimates for each mask to evaluate its quality. In medical image analysis, automated quality control is critical to ensure the reliability and safety of the deep model outputs (Kohlberger et al., 2012). For example, Wang et al. (2020d) employed deep generative models for learnable quality control in cardiac MRI segmentation, where the predicted Dice scores showed a strong linear relationship with the real ones. Additionally, Billot et al. (2023) used an additional neural network to predict the Dice coefficient of brain tissue segmentation results. Overall, learnable performance estimation can accurately predict the quality of model outputs. Hence, delving deeper into their potential for uncertainty-based AL is crucial to effectively tackle the issue of over-confidence.
@@
Among all the mentioned methods in §3.1, adversarial-based uncertainty currently has limited applications in AL of medical image analysis. Since the adversarial samples tend to be close to the classification boundary, they can be regarded as uncertain samples, and selecting them for training can potentially improve the trained model’s robustness. Exploring such ideas in medical image analysis, especially in the federated learning scenario, could be an interesting topic for future work.
@@ 7.2 Towards Active Learning with Better Representativeness
Representativeness-based AL effectively utilizes feature representations and data distributions for sample selection. Cover-based and discrepancy-based AL methods implicitly capture the data distribution, whereas density-based AL explicitly estimates it. However, the latter requires supplementary strategies to ensure diversity. For discrepancy-based AL, we can opt for a better metric of the distance between two probability distributions (Zhao et al., 2022). Besides, discrepancy-based AL has limited applications in medical image analysis currently. Finding a proper metric for medical images considering their special characteristics could be a promising direction for the future development of AL in the medical imaging domain.
As the core of density-based AL, density estimation in high-dimensional spaces has always been challenging. Popular density estimation methods, such as kernel density estimation and GMM, can encounter challenges when applied in high-dimensional spaces. In future research, we can consider introducing density estimators tailored to high-dimensional spaces. Advanced tools like normalizing flow (Papamakarios et al., 2021) could be an appropriate choice in density estimation in high-dimensional spaces.
@@ 7.3 Towards Active Learning with Weak Annotation
In §4.3, we discuss region-based active learning, which only requires region-level annotation of a sample. However, annotating all pixels within the region is still needed. Several existing works have incorporated weak annotations with AL to simplify the task for annotators. In object detection tasks, Vo et al. (2022) trained deep models with image-level annotation. They selected samples with box-in-box prediction results and annotated them with bounding boxes. Moreover, Lyu et al. (2023) adopted disagreement to choose which objects are worth annotating. Rather than annotating all objects within the image, they only required box-level annotations for a subset of objects. In AL of instance segmentation, Tang et al. (2022a) only required annotations for each object’s class label and bounding box, without the annotation of fine-grained segmentation masks. In future research, AL based on weak annotations is a direction worthy of in-depth exploration.
@@ 7.4 Towards Active Learning with Better Generative Models
In §4.4, we summarize the applications of generative models in AL. However, existing works have mainly focused on using GANs as sample generators. Recently, diffusion models (Kazerouni et al., 2023) have advanced in achieving state-of-the-art generative quality. Furthermore, text-to-image diffusion models, represented by Stable Diffusion (Rombach et al., 2022), have revolutionized the image generation domain. Their high-quality, text-guided generation results enable a more flexible image generation. With the use of ControlNet Zhang et al. (2023a), the diffusion models could learn to follow a more detailed condition like a sketch or segmentation mask. Exploring the potential of diffusion models in deep AL is a promising avenue for future research.
@@ 7.5 Towards Active Learning with Foundation Models
With the rise of visual foundational models, such as contrastive language-image pre-training (CLIP) (Radford et al., 2021) and SAM (Kirillov et al., 2023), and large language models (LLMs) like GPT-4 (OpenAI, 2023), deep learning in medical image analysis and computer vision is undergoing a paradigm shift. These foundational models (Bommasani et al., 2021) offer new opportunities for the development of AL.
AL is closely related to the training paradigms in deep learning of computer vision and medical image analysis. From the initial approach of train-from-scratch to the “pre-train-finetune” strategy using supervised or self-supervised pre-trained models, these paradigms usually require fine-tuning the entire network. Foundation models contain a wealth of knowledge. When combined with recently emerging parameter-efficient fine tuning (PEFT) or prompt tuning techniques (Hu et al., 2021; Jia et al., 2022), we can tune only a minimal subset of model weights (e.g., 5%) for rapid transfer to downstream tasks. As the number of fine-tuned parameters decreases, AL has the potential to further reduce the number of required annotated samples. Bai et al. (2023) integrated prompt tuning with AL in liver tumor segmentation. A segmentation model trained on publicly available datasets was transferred to in-house datasets via a novel prompt updater. With a mixed AL strategy of uncertainty and diversity, the proposed method reached the comparable performance of fully supervised tuning using around 5% of samples and 6% of tunable parameters. Therefore, it is essential to investigate the applicability of existing AL under PEFT or prompt tuning and explore the most suitable AL strategies for PEFT.